Circular No. 7174

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

1SAX J0835.9+5118 G. Gandolfi, A. Coletta, and D. Ricci, BeppoSAX Scientific Operation

Center, Telespazio, Rome; J. Heise and J. in’t Zand, Space Research Organization of the Netherlands, Utrecht; and L. Piro, E. Costa, and M. Feroci, Istituto di Astrofisica Spaziale, CNR, Rome, write: “On May 20.08539 UT a fast x-ray transient, which we designate 1SAX J0835.9+5118, has been detected in BeppoSAX Wide Field Camera unit 1 (2–28 keV). The event lasted ∼ 10 s, with a peak flux of 1 Crab in the range 2–25 keV. The burst shows two peaks; in coincidence with the second, harder burst, the GRB Monitor (40–700 keV) obtained a marginal detection (4σ signal) corresponding to a flux of ∼ 1.8 Crab. The position is α = 8^{h}35^{m}56^{s}, δ = +51^{o}18′.6 (equinox 2000.0), with an error radius of 3′. No catalogued source at any wavelength lies in the error box. Furthermore, no quiescent emission was found from this source during 3 hr following the outburst, with an upper limit of 6 mCrab (2–10 keV), and in a combination of 1996–1999 data, with an upper limit of 0.6 mCrab. The normalized flux differences F_{1} − F_{2} and F_{2} − F_{3} (where F_{1}, F_{2}, and F_{3} are the fluxes in the bands 20–26, 7–20, and 3–7 keV) are −0.62 and +0.05, respectively. These values are typically softer than GRBs and fall in the same range as type-I x-ray bursters. Follow-up observations are encouraged to investigate the nature of this object.”

COMET C/1998 T1 (LINEAR) Visual m_{1} estimates: 1998 Nov. 19.80 UT, 13.8 (M. Lehký, Hradec

Králové, Czech Republic, 0.42-m reflector); Dec. 17.79, 13.6 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 1999 Jan. 12.76, 13.1 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); May 19.61, 11.5 (N. Biver, Oahu, HI, 0.26-m reflector).

COMET C/1999 H3 (LINEAR) Visual m_{1} estimates: May 5.88 UT, 13.8 (Hornoch); 9.89, 13.5 (Lehký);

16.07, 13.7 (J. Carvajal, Ciudad Real, Spain, 0.45-m reflector); 17.92, 13.1 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector); 18.96, 13.3 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 19.90, 13.6 (Lehký).

1999 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green